# Coordination Requires Simplification: Thermodynamic Bounds on Multi-Objective Compromise in Natural and Artificial Intelligence

### v1

@@ Abstract
Information-processing systems coordinating across multiple agents and objectives face fundamental thermodynamic constraints. We show that solutions with maximum utility to act as coordination focal points have much higher selection pressure for being findable across agents rather than accuracy. We derive that the information-theoretic minimum description length of coordination protocols to precision $\varepsilon$ scales as $L(P)\geq NK\log_{2}K+N^{2}d^{2}\log(1/\varepsilon)$ for $N$ agents with $d$ potentially conflicting objectives and internal model complexity $K$. This scaling forces progressive simplification, with coordination dynamics changing the environment itself and shifting optimization across hierarchical levels. Moving from established focal points requires re-coordination, creating persistent metastable states and hysteresis until significant environmental shifts trigger phase transitions through spontaneous symmetry breaking. We operationally define coordination temperature to predict critical phenomena and estimate coordination work costs, identifying measurable signatures across systems from neural networks to restaurant bills to bureaucracies. Extending the topological version of Arrow’s theorem on the impossibility of consistent preference aggregation, we find it recursively binds whenever preferences are combined. This potentially explains the indefinite cycling in multi-objective gradient descent and alignment faking in Large Language Models trained with reinforcement learning with human feedback. We term this framework Thermodynamic Coordination Theory (TCT), which demonstrates that coordination requires radical information loss.
I Introduction
@@
Agents in complex systems face fundamental information-theoretic constraints as finite resources must model environments of arbitrary complexity. Working memory limitations force agents to function with a smaller subset of available information, creating a bottleneck [1]. While coordination with other agents having different objectives can increase collective capacity, information is both compressed and lost at every emergent hierarchical level [2]. This principle that any tractable macroscopic description requires information loss by coarse-graining over microscopic degrees of freedom, has deep parallels in the quantum origins of the classical world [3]. We elaborate how this establishes a fundamental tension between scaling and accuracy. The Free Energy Principle (FEP) describes how individual systems minimize surprise by updating internal models using prediction errors from sensory inputs through Bayesian inference [4, 5], but no current formalisms address the information-theoretic costs and emergent behavior when multiple agents must coordinate their predictive models across potentially conflicting objectives. This is the gap this paper addresses.
@@
Anderson showed ‘more is different’ [6], i.e., collective properties emerge at scale that cannot be reduced to constituent behaviors. We quantify this emergence mechanism for information-processing systems: coordination requires radical simplification because agents must converge on Schelling focal points [7] despite having different objectives and limited communication capacity. This convergence forces systematic information loss. When selection pressures for findability, accuracy, and other objectives compete, emergent dynamics determine which information the system preserves and which it discards.
@@
This framework builds upon computational thermodynamics results to establish our scaling bounds. Landauer’s principle provides the minimum energy for information erasure [8]. Bennett extended this thermodynamic cost to general computation [9]. Wolpert’s synthesis connects stochastic thermodynamics to computation theory [10], while his no free lunch theorems [11] and work on physical limits of inference [12] establish that no universal optimization exists. We leverage these results to prove multi-objective coordination cannot have a general solution and that coordination has thermodynamic costs that scale super-linearly. In parallel, information-theoretic approaches to prediction [13] and decision-making [14] establish optimal behavior and energetic costs that we will show multiply when agents must coordinate. Information theory [15] provides the mathematical foundation for our quantification of these costs.
@@
This paper synthesizes insights from multiple fields but is largely self-contained. While we cite literature from statistical mechanics, information theory, computer science, and social choice theory for context, the main arguments require only basic familiarity with information theory (measured in bits) and probability. Readers from different backgrounds may find some sections more intuitive than others. The rest of the paper is organized as follows. In Sec. II, we illustrate our theoretical framework with the example of bill-splitting or “bistromathics”. The main results and their derivations are provided in Sec. III. The expected critical phenomena and physical parameters are described in Sec. IV. We conclude and discuss the implications of TCT in Sec. V.
II Restaurant Bill Splitting as Illustration
@@
For $N=4$ diners to completely coordinate with $d=2$ (items and fairness): Each diner must communicate items consumed, e.g. “I had the soup (\$8), half the calamari (\$6), two beers (\$12)” yields $K\approx 20$ bits (prices, quantities, shares, customizations), and $\binom{4}{2}\times\binom{2}{2}=6$ pairwise reconciliation ($\sim$ 5 bits each), e.g. “Did we split the calamari evenly?”. Total communication: 80 + 30 = 110 bits. For an explanation of multiplication in the pairwise reconciliation, see the topological obstruction created by preference cycling of objectives described in the topological derivation of Theorem 1 in Supplementary Section I.
@@
This problem termed “bistromathics” by Douglas Adams in book two of his five-part Hitchhiker’s Guide to the Galaxy trilogy shows bill calculations become non-absolute (focal) points [30], representing phase transition to simplified coordination. Conflict resolution protocols include restaurants imposing terms like “3 cards maximum” or “20 % minimum gratuity for parties of 6 or more”, and agents declaring they do not use digital peer-to-peer payments or carry cash. Varying $N$ is qualitatively explored in Section III of the Supplementary. Finding the non-absolute focal points (Adams’ recipriversexclusions) involves Level II chaos [31] which is beyond this paper’s scope.
Level II Chaos (Definition): A term for a specific subclass of Complex Adaptive Systems (CAS), the formal framework for systems of interacting, adaptive agents [32]. In these dynamics, agents continuously attempt to update their internal models to predict the system’s future state, which in turn alters their and the system’s behavior. This reflexive feedback loop is particularly evident in financial markets, where the collective forecasting efforts of participants shape the very price movements they attempt to predict [33]. Unlike Level I chaos (e.g., weather, N-body dynamics), whose dynamics are unaffected by prediction, Level II chaos is an emergent property of this process: the unpredictability of macro-state is compounded by the act of trying to forecast it. While Level II systems exhibit their own reflexive dynamics, they remain susceptible to Level I influences, for instance, stock returns correlating with local sunshine [34].
This definition is crucial for our framework as coordination attempts fundamentally alter the landscape being coordinated. Established focal points exhibit metastability and path-dependent hysteresis, making re-coordination thermodynamically costly. Level II Chaos with the Bistromathics example is illustrated in Section III of the Supplementary.
III Mathematical Derivations
@@
III.1 Theorem 1: Coordination Info Scale as $N^{2}d^{2}$
@@
$L(P)$ $\geq L_{\rm{models}}+L_{\rm{comm}}$
$=N\bar{K}\log\bar{K}\,h(\rho)+\binom{N}{2}\frac{d(d+3)}{2}\log(1/\varepsilon)$ (1)
@@
In practice, model complexity $K$ may depend on both $N$ and $d$. Each additional objective typically requires additional state representation, suggesting K scales at least linearly with $d$ [36], i.e. $K\geq K_{b}+K_{\delta}d$ for some base complexity $K_{b}$ and $K_{\delta}$ per objective. The Bistromathics example in Section II uses $K_{b}=0,\ K_{\delta}=10$. Additionally, agents may need to model other agents’ states, implying $K$ could scale with $N$ for small groups. However, cognitive limits in humans [1], and bounded rationality forces categorical simplified representations rather than individual models at large $N$ [37]. Here we treat $K$ as largely constant to derive lower bounds, noting that any dependence $K(N,d)$ would only worsen the coordination challenge.
@@
$L_{\rm{comm}}$ $\geq\frac{N(N-1)}{2}\cdot I_{pair}$
$=\binom{N}{2}\frac{d(d+3)}{2}\log(1/\varepsilon)\cdot\{1+\mathcal{O}(1/d)\}$ (4)
@@
We provide an independent derivation of Theorem 1 via the topology of preference spaces from the Chichilnisky extension of Arrow’s Impossibility theorem [21], confirming the $\rm{N^{2}d^{2}}$ scaling through geometric reasoning (Supplementary Section I). This convergence of information-theoretic and topological approaches suggests the scaling is fundamental, not algorithmic. Therefore, no protocol innovation can reduce this bound without sacrificing coordination accuracy. Figure 1 panel (a) shows a graphical representation of how $L(P)$ scales with $N$ and $d$.
@@
where the first term represents model sharing, the second within-group coordination, and the third between-group coordination. See Supplementary Section II for full derivations. Minimizing in the limit of large $N$ over $M$ yields $M_{opt}=N^{2/3}$ and:
@@
$L_{\rm{multi}}$ $\approx N\bar{K}\log\bar{K}\,h(\rho)\frac{b^{2}}{b-1}$
$+\frac{1}{4}Nb^{2}d_{\rm{H}}(d_{\rm{H}}+3)\log(1/\varepsilon)$ (9)
@@
III.2 Theorem 2: Findability Dominates Accuracy
@@
$U(s)=\Omega[A(s)]\times P(\text{coordinate on }s)$ (11)
where $\Omega:[0,1]\rightarrow\mathbb{R}^{+}$ is an unspecified monotonic function weighting accuracy. This multiplicative form follows from network reliability theory [40] where system utility requires both solution quality and successful propagation through the agent network, analogous to series reliability or bootstrap percolation [41]. Note that this formulation assumes independent discovery and acceptance probabilities, though correlations could be incorporated.
@@
$\frac{\partial U}{\partial F_{i}}=\Omega[A(s)]\times\prod_{j\neq i}F_{j}(s)\sim\Omega[A(s)]\times P(s)$ (12)
@@
$\frac{\partial U}{\partial A}=\Omega^{\prime}[A(s)]\times P(\text{coordinate on }s)$ (13)
@@
$\frac{\partial U/\partial F_{i}}{\partial U/\partial A}=\frac{\Omega[A(s)]}{F_{i}(s)\times\Omega^{\prime}[A(s)]}\gg 1$ (14)
@@
IV Critical Phenomena, and Phase Transitions
We define measurable parameters expected phenomenologically from our two theorems. The definitions are subject to revisions, and re-derivation from a microscopic Hamiltonian in subsequent work. Empirical measurement presents challenges discussed in the next Section.
IV.1 Coordination Temperature
@@
IV.2 Renormalization Group Flow
@@ IV.3 Critical Transitions
The $\log(N)$ scaling emerges from hierarchical coordination (Equation III.1) requiring $\mathcal{O}\{\log(N)\}$ levels, each introducing noise. This is analogous to how network diameter scales as $\log(N)$ in small-world networks. The RG flow becomes unstable when thermal fluctuations exceed the per-agent capacity to maintain coordination, yielding the critical temperature. Systems exhibit metastability, persisting at suboptimal focal points until fluctuations overcome coordination barriers. Beyond $T_{c,co}$, accurate coordination becomes thermodynamically prohibitive. The order parameter $\psi=(K_{absolute}-K_{relative})/K_{absolute}$ measuring model deviation should exhibit rapid transitions near $T_{c,co}$, with width $\Delta T\sim\sqrt{T_{c,co}}$. Here, $K_{absolute}$ is the true system complexity, and $K_{relative}$ is the coordinated model complexity, with $K_{relative}\approx K_{0}$ at the focal point. This order parameter displays hysteresis characteristic of first-order transitions with groups maintaining $\psi$ values despite crossing $T_{c,co}$ until perturbations trigger reorganization.
@@
IV.4 Energy for Metastable Coordination
From Theorem 2, we expect coordination across larger groups and more objectives requires simpler coordination focal points. However, since each hierarchical level (Equation III.1) introduces agent model prediction errors due to information loss, complexity cannot spontaneously increase or be maintained at a simpler level without external work.

### v3

@@ Abstract
Information-processing systems that coordinate multiple agents and objectives face fundamental thermodynamic constraints. We show that solutions with maximum utility to act as coordination focal points have a much higher selection pressure for being findable across agents rather than accuracy. We derive that the information-theoretic minimum description length of coordination protocols to precision $\varepsilon$ scales as $L(P)\geq NK\log_{2}K+N^{2}d^{2}\log(1/\varepsilon)$ for $N$ agents with $d$ potentially conflicting objectives and internal model complexity $K$. This scaling forces progressive simplification, with coordination dynamics changing the environment itself and shifting optimization across hierarchical levels. Moving from established focal points requires re-coordination, creating persistent metastable states and hysteresis until significant environmental shifts trigger phase transitions through spontaneous symmetry breaking. We operationally define coordination temperature to predict critical phenomena and estimate coordination work costs, identifying measurable signatures across systems from neural networks to restaurant bills to bureaucracies. Extending the topological version of Arrow’s theorem on the impossibility of consistent preference aggregation, we find it recursively binds whenever preferences are combined. This potentially explains the indefinite cycling in multi-objective gradient descent and alignment faking in Large Language Models trained with reinforcement learning with human feedback. We term this framework Thermodynamic Coordination Theory (TCT), which demonstrates that coordination requires radical information loss.
keywords:
Scaling laws of complex systems, Thermodynamics of Computation, Information Theory, Statistical Mechanics, Chaos & nonlinear dynamics, Artificial Intelligence, Social Physics
articletype: Paper
1 Introduction
@@
Agents in complex systems face fundamental information-theoretic constraints as finite resources must model environments of arbitrary complexity. Working memory limitations force agents to function with a smaller subset of available information, thereby creating a bottleneck [1]. While coordination with other agents with different objectives can increase collective capacity, information is both compressed and lost at every emergent hierarchical level [2]. This principle that any tractable macroscopic description requires information loss by coarse-graining over microscopic degrees of freedom, has deep parallels in the quantum origins of the classical world [3]. We elaborate how this establishes a fundamental tension between scaling and accuracy. The Free Energy Principle (FEP) describes how individual systems minimize surprise by updating internal models using prediction errors from sensory inputs through Bayesian inference [4, 5], but no current formalisms address the information-theoretic costs and emergent behavior when multiple agents must coordinate their predictive models across potentially conflicting objectives. This is the gap this paper addresses.
@@
Anderson showed ‘more is different’ [6], i.e., collective properties emerge at scale that cannot be reduced to constituent behaviors. We quantify this emergence mechanism for information-processing systems: coordination requires radical simplification because agents must converge on Schelling focal points [7] despite having conflicting objectives and limited communication capacity. This convergence forces systematic information loss. When selection pressures for findability, accuracy, and other objectives compete, emergent dynamics determine which information the system preserves and discards.
@@
This framework builds upon computational thermodynamic results to establish our scaling bounds. Landauer’s principle provides the minimum energy required for information erasure [8]. Bennett extended this thermodynamic cost to general computation [9]. Wolpert’s synthesis connects stochastic thermodynamics to computation theory [10], while the no free lunch theorems [11] and works on physical limits of inference [12] establish that no universal optimization exists. We leverage these results to prove multi-objective coordination cannot have a general solution and that coordination has thermodynamic costs that scale super-linearly. In parallel, information-theoretic approaches to prediction [13] and decision-making [14] establish optimal behavior and energetic costs that we will show multiply when agents must coordinate. Information theory [15] provides the mathematical foundation for our quantification of these costs.
@@
This paper synthesizes insights from multiple fields but is largely self-contained. While we cite literature from statistical mechanics, information theory, computer science, and social choice theory for context, the main arguments require only basic familiarity with information theory (measured in bits) and probability. Readers from different backgrounds may find some sections more intuitive than others. The rest of the paper is organized as follows. In Sec. 2, we illustrate our theoretical framework with the example of bill-splitting or “bistromathics”. The main results and their derivations are provided in Sec. 3. The expected critical phenomena and physical parameters are described in Sec. 4. We conclude and discuss the implications of TCT in Sec. 5.
2 Restaurant Bill Splitting as Illustration
@@
For $N=4$ diners to completely coordinate with $d=2$ (items and fairness): Each diner must communicate items consumed, e.g. “I had the soup (\$8), half the calamari (\$6), two beers (\$12)” yields $K\approx 20$ bits (prices, quantities, shares, customizations), and $\binom{4}{2}\times\binom{2}{2}=6$ pairwise reconciliation ($\sim$ 5 bits each), e.g. “Did we split the calamari evenly?”. Total communication: 80 + 30 = 110 bits. For an explanation of multiplication in the pairwise reconciliation, see the topological obstruction created by preference cycling of objectives described in the topological derivation of Theorem 1 in Supplementary Section 1.
@@
This problem termed “bistromathics” by Douglas Adams in book two of his five-part Hitchhiker’s Guide to the Galaxy trilogy shows bill calculations become non-absolute (focal) points [30], representing phase transition to simplified coordination. Conflict resolution protocols include restaurants imposing terms like “3 cards maximum” or “20 % minimum gratuity for parties of 6 or more”, and agents declaring they do not use digital peer-to-peer payments or carry cash. Varying $N$ is qualitatively explored in Section 3 of the Supplementary. Finding the non-absolute focal points (Adams’ recipriversexclusions) involves Level II chaos [31] which is beyond this paper’s scope.
Level II Chaos (Definition): A term for a specific subclass of Complex Adaptive Systems (CAS), the formal framework for systems of interacting, adaptive agents [32]. Level II chaos describes systems whose trajectories are unknowable because of the necessary recall and loss of information required to transition between states. The subset of information recalled is path-dependent and unknowable a priori, as agents must select which information to preserve when complete description length exceeds system capacity. In these dynamics, agents continuously attempt to update their internal models to predict the system’s future state, which in turn alters their and the system’s behavior. This reflexive feedback loop is particularly evident in financial markets, where the collective forecasting efforts of participants shape the very price movements they attempt to predict [33]. Unlike Level I chaos (e.g., weather, N-body dynamics), whose dynamics are unaffected by prediction, Level II chaos is an emergent property of this process: the unpredictability of macro-state is compounded by the act of trying to forecast it as it changes accessible information. While Level II systems exhibit their own reflexive dynamics, they remain susceptible to Level I influences, for instance, stock returns correlating with local sunshine [34].
This definition is crucial for our framework as coordination attempts fundamentally alter the landscape being coordinated. Established focal points exhibit metastability and path-dependent hysteresis, making re-coordination thermodynamically costly. While this characterization provides a useful conceptual framework for understanding reflexive dynamics in complex systems, mathematical formalization and empirical validation remain topics for future work. Level II Chaos with the Bistromathics example is illustrated in Section 3 of the Supplementary.
3 Mathematical Derivations
@@
3.1 Theorem 1: Coordination Info Scale as $N^{2}d^{2}$
@@
$L(P)$ $\geq L_{\rm{models}}+L_{\rm{comm}}$ (1)
$=N\bar{K}\log\bar{K}\,h(\rho)+\binom{N}{2}\frac{d(d+3)}{2}\log(1/\varepsilon)$
@@
In practice, model complexity $K$ may depend on both $N$ and $d$. Each additional objective typically requires additional state representation, suggesting K scales at least linearly with $d$ [36], i.e. $K\geq K_{b}+K_{\delta}d$ for some base complexity $K_{b}$ and $K_{\delta}$ per objective. The Bistromathics example in Section 2 uses $K_{b}=0,\ K_{\delta}=10$. Additionally, agents may need to model other agents’ states, implying $K$ could scale with $N$ for small groups. However, cognitive limits in humans [1], and bounded rationality forces categorical simplified representations rather than individual models at large $N$ [37]. Here we treat $K$ as largely constant to derive lower bounds, noting that any dependence $K(N,d)$ would only worsen the coordination challenge.
@@
$L_{\rm{comm}}$ $\geq\frac{N(N-1)}{2}\cdot I_{pair}$ (4)
$=\binom{N}{2}\frac{d(d+3)}{2}\log(1/\varepsilon)\cdot\{1+\mathcal{O}(1/d)\}$
@@
We provide an independent derivation of Theorem 1 via the topology of preference spaces from the Chichilnisky extension of Arrow’s Impossibility theorem [21], confirming the $\rm{N^{2}d^{2}}$ scaling through geometric reasoning (Supplementary Section 1). This convergence of information-theoretic and topological approaches suggests the scaling is fundamental, not algorithmic. Therefore, no protocol innovation can reduce this bound without sacrificing coordination accuracy. Figure 1 panel (a) shows a graphical representation of how $L(P)$ scales with $N$ and $d$.
@@
where the first term represents model sharing, the second within-group coordination, and the third between-group coordination. See Supplementary Section 2 for full derivations. Minimizing in the limit of large $N$ over $M$ yields $M_{opt}=N^{2/3}$ and:
@@
$L_{\rm{multi}}$ $\approx N\bar{K}\log\bar{K}\,h(\rho)\frac{b^{2}}{b-1}$ (9)
$+\frac{1}{4}Nb^{2}d_{\rm{H}}(d_{\rm{H}}+3)\log(1/\varepsilon)$
@@
3.2 Theorem 2: Findability Dominates Accuracy
@@
$U(s)=\Omega[A(s),M,N]\cdot P(\text{coordinate on }s)$ (11)
where $\Omega:\{[0,1],\mathbb{N},\mathbb{N}\}\rightarrow\mathbb{R}^{+}$ is an unspecified monotonic function of $A$ weighting accuracy and agent thresholds. This multiplicative form follows from network reliability theory [40] where system utility requires both solution quality and successful propagation through the agent network, analogous to series reliability or bootstrap percolation [41]. Note that this formulation assumes independent discovery and acceptance probabilities, though correlations could be incorporated. While other utility functions like additive or threshold models exist, they might be less suited for describing systems where emergence and collective action are paramount. In such systems, a solution with zero findability or veto power must have zero system-wide utility, a condition that the multiplicative form with stakeholder thresholds naturally captures.
@@
$\frac{\partial U}{\partial F_{i}}=\Omega[A(s)]\cdot\prod_{j\neq i}F_{j}(s);\ \lim_{F_{i}\to 1}\frac{\partial U}{\partial F_{i}}=\Omega\cdot P(s)$ (12)
@@
$\frac{\partial U}{\partial A}=\Omega^{\prime}[A(s)]\cdot P(\text{coordinate on }s)$ (13)
@@
$\frac{\partial U/\partial F_{i}}{\partial U/\partial A}=\frac{\Omega[A(s)]}{F_{i}(s)\cdot\Omega^{\prime}[A(s)]}\gg 1$ (14)
@@
4 Critical Phenomena, and Phase Transitions
We define measurable parameters expected phenomenologically from our two theorems. The definitions are subject to revisions, and re-derivation from a microscopic Hamiltonian in subsequent work. Empirical measurement presents challenges discussed in Section 5.
4.1 Coordination Temperature
@@
4.2 Renormalization Group Flow
@@ 4.3 Critical Transitions
The $\log(N)$ scaling emerges from hierarchical coordination (Equation 9) requiring $\mathcal{O}\{\log(N)\}$ levels, each introducing noise. This is analogous to how network diameter scales as $\log(N)$ in small-world networks. The RG flow becomes unstable when thermal fluctuations exceed the per-agent capacity to maintain coordination, yielding the critical temperature. Systems exhibit metastability, persisting at suboptimal focal points until fluctuations overcome coordination barriers. The focal points in this framework are analogous to priors in predictive processing and Bayesian brain models [25, 4]. Both represent informational attractors that minimize surprise through compression.
Beyond $T_{c,co}$, accurate coordination becomes thermodynamically prohibitive. The order parameter $\psi=(K_{absolute}-K_{relative})/K_{absolute}$ measuring model deviation should exhibit rapid transitions near $T_{c,co}$, with width $\Delta T\sim\sqrt{T_{c,co}}$. Here, $K_{absolute}$ is the true system complexity, and $K_{relative}$ is the coordinated model complexity, with $K_{relative}\approx K_{0}$ at the focal point. This order parameter displays hysteresis characteristic of first-order transitions with groups maintaining $\psi$ values despite crossing $T_{c,co}$ until perturbations trigger reorganization.
@@
4.4 Energy for Metastable Coordination
From Theorem 2, we expect coordination across larger groups and more objectives requires simpler coordination focal points. However, since each hierarchical level (Equation 9) introduces agent model prediction errors due to information loss, complexity cannot spontaneously increase or be maintained at a simpler level without external work.